For the following reaction:

2NOCl(g) → 2NO(g) + Cl2(g)

     determine:

     a) the differential rate law

     b) the integrated rate law

c) the value for the rate constant

d) the concentration of NOCl(g) remaining after 275 s in each trial

given:

Trial

[NOCl]0 (mol/L)

Initial Rate (mol/L•s)

1

3.15

7.88 × 10–3

2

6.30

1.57 × 10–2

3

12.6

3.15 × 10–2
Answer #1

- rale = a let the differential rate law will be a Pove = K[noci]. Now rato a when [roci] = 3:15, r= 788x10-3 sv 7.88810-3 =Nowy ra 7189 x10-3 [nvei] = 3.15 Ké /{nect] 7.88810-3 mothest - 2 3:15 boll = 2.5 X10-35-